OpenStudy (thomas5267):

Ask for my friend, I don't know this either. In how many ways can 5 letters be mailed if there are: a) 2 mailboxes available b) 4 mailboxes available Edit: I erroneously wrote that 5 mailboxes is available in b).

OpenStudy (thomas5267):

My intuition is that the answer to the first one is 6 as: A B 0 5 1 4 2 3 3 2 4 1 5 0 But the answer is 25. Apparently the letters are distinguishable.

OpenStudy (thomas5267):

@Kainui HELP!!!!!!

OpenStudy (rational):

If the letters are distinguishable then each letter has 2 possible states : {box1, box2} So total possible states for sending 5 letters would be : 2*2*2*2*2 = 2^5

OpenStudy (phi):

if the letters are distinguishable I get 32 ways, not 25

OpenStudy (rational):

thats for part a

OpenStudy (thomas5267):

Me too. I am so confused right now.

OpenStudy (thomas5267):

The answer to part a is 25 and part b is 625 = 25*25.

OpenStudy (thomas5267):

The book could well be wrong. It happened before but I need a solid answer to say the book is wrong.

OpenStudy (thomas5267):

Is it possible that the mail is indistinguishable but some mail can be lost in post? LOL!

OpenStudy (rational):

list them all if that states arguement doesn't look solid

OpenStudy (rational):

call the letters : l1, l2, l3, l4, l5 and boxes : b1, b2

OpenStudy (thomas5267):

It is basically \[ \sum_{k=0}^5\binom{5}{k}=2^5 \]

OpenStudy (rational):

thats the total number of ways of choosing 0->5 letters for first box second box is determined after that so yeah

OpenStudy (thomas5267):

How do you deal with the second part though? 4^5?

OpenStudy (rational):

Yes assuming the letters are distinguishable

OpenStudy (rational):

If they're not distinguishable then the answer is simply 6 for part a and 56 for part b

OpenStudy (phi):

oh, it looks like the answer is doing (# of letters)^(# of boxes) so 5^2 = 25 and 5^4= 625

OpenStudy (thomas5267):

The book is dumb? \[ 2^5=32\\ 5^2=25\quad\text{book's answer to part a}\\ 4^5=1024\\ 5^4=625\quad\text{book's answer to part b} \]

OpenStudy (phi):

so the book is pretty close

OpenStudy (rational):

the author must be drunk when writing solution

OpenStudy (thomas5267):

*facepalm* I wouldn't call that pretty close. Yes the author must be drunk.

OpenStudy (thomas5267):

It is beyond words!

OpenStudy (phi):

they used both numbers, and an exponent.

OpenStudy (rational):

whats the complete question anyways

OpenStudy (thomas5267):

That is the exact wording of that question LOL!

OpenStudy (thomas5267):

I know the question is worded poorly and it is pretty common in this textbook.

OpenStudy (rational):

im just wondering if there is a way to get to textbook answer by putting constraints like 1) making letters indistinguishable 2) each box must have atleast 1 letter etc..

OpenStudy (thomas5267):

If those constraints are true, then part a would be 4 not 25.